OpenStudy (anonymous):

The function F(x)=e^(-3x^2) has 2 inflection points. Find the x-coordinate of the inflection point when x>0, correct to 2 decimal places.

OpenStudy (anonymous):

King, this is the same problem TheFigure was working on.... he just couldn't figure it out. (pun not intended lol)

OpenStudy (anonymous):

@amistre64

OpenStudy (kinggeorge):

The method he was using is definitely what you want to use. He probably just took one of the derivatives wrong.

OpenStudy (anonymous):

could u go see where he went wrong

OpenStudy (kinggeorge):

Can you check if a certain answer is correct, and if so I'll tell you how to do it?

OpenStudy (kinggeorge):

Try\[{1 \over \sqrt{6} }\approx 0.41\]

OpenStudy (anonymous):

thats correct!!

OpenStudy (kinggeorge):

After looking through Figure's work, I'm pretty sure he took the first derivative wrong. When he wrote\[f'(x)= -6x^2e^{-3x^2}\]It should have been\[f'(x)=-6xe^{-3x^2}\]Since the derivative of \(-3x^2\) is \(-6x\). Now, if we use this new \(f'\), we have that \[f''(x)=6e^{-3x^2}(6x^2-1)\]Thus, \(6x^2-1=0\) which means that \(x=\pm1/\sqrt{6}\) But since we're interested in positive x-values, \(x=1/\sqrt{6}\)

OpenStudy (anonymous):

ok thanks for your time man. much appreciated.

OpenStudy (kinggeorge):

You're welcome. Once again, sorry I wasn't able to answer this sooner.

OpenStudy (anonymous):

at least you helped. thats all that matters.